OpenStudy (anonymous):

Determine the electronegativity between the atoms of each molecule of H2O, NH3, CH4, and CO2

OpenStudy (btaylor):

H2O: 2 O-H bonds, 3.5 - 2.1 = 1.4 NH3: 3 N-H bonds, 3.0 - 2.1 = 0.9 CO2: 2 C-O bonds, 3.5 - 2.5 = 1.0